# EditReward: A Human-Aligned Reward Model for Instruction-Guided Image Editing

### v1

@@ Abstract
Recently, we have witnessed great progress in image editing with natural language instructions. Several closed-source models like GPT-Image-1, Seedream, and Google-Nano-Banana have shown highly promising progress. However, the open-source models are still lagging. The main bottleneck is the lack of a reliable reward model to scale up high-quality synthetic training data. To address this critical bottleneck, we built EditReward, trained with our new large-scale human preference dataset, meticulously annotated by trained experts following a rigorous protocol containing over 200K preference pairs. EditReward demonstrates superior alignment with human preferences in instruction-guided image editing tasks. Experiments show that EditReward achieves state-of-the-art human correlation on established benchmarks such as GenAI-Bench, AURORA-Bench, ImagenHub, and our new EditReward-Bench, outperforming a wide range of VLM-as-judge models. Furthermore, we use EditReward to select a high-quality subset from the existing noisy ShareGPT-4o-Image dataset. We train Step1X-Edit on the selected subset, which shows significant improvement over training on the full set. This demonstrates EditReward’s ability to serve as a reward model to scale up high-quality training data for image editing. Furthermore, its strong alignment suggests potential for advanced applications like reinforcement learning-based post-training and test-time scaling of image editing models. EditReward with its training dataset will be released to help the community build more high-quality image editing training datasets.
@@ 1 Introduction
Figure 2: Statistics of our EditReward-Data and EditReward-Bench.
@@ 2.1 The EditReward-Data Construction
EditReward-Data contains 9557 instruction–image pairs collected from six established editing benchmarks: GEdit-Bench (606) (Liu et al., 2025b), ImgEdit-Bench (737) (Ye et al., 2025), MagicBrush (1,053) (Zhang et al., 2024a), AnyEdit (1,250) (Yu et al., 2025), EmuEdit (5,611) (Sheynin et al., 2024), and an internal set (300). This aggregation ensures broad coverage of semantically grounded and executable editing instructions. For each instruction, we generated 12 candidate images using six state-of-the-art models: Step1X-Edit (Liu et al., 2025b), Flux-Kontext (BlackForestLabs et al., 2025), Qwen-Image-Edit (Wu et al., 2025a), BAGEL (Deng et al., 2025), Ovis-U1 (Wang et al., 2025a), and OmniGen2 (Wu et al., 2025b), with multiple random seeds to avoid model bias. Seven candidates were randomly sampled for human evaluation. Annotators scored each image on a 4-point Likert scale (1 = Poor and 4 = Excellent) along two dimensions: Instruction Following (semantic accuracy, completeness, and no unprompted changes) and Visual Quality (plausibility, artifact-free rendering, and aesthetics). This rubric yields more informative labels than single-score schemes. Details of the annotation protocol and quality-control process are provided in the Appendix  A.2. Comprehensive statistics of the dataset are provided in Table 1 and Figure 2. EditReward-Data is unique in combining large-scale, expert human annotation and a multi-dimensional scoring rubric, making it a strong foundation for training editing reward models.
@@ 3.2 Multi-Dimensional Uncertainty-Aware Ranking
This strategy forces the model to reconcile seemingly contradictory signals for the same input pair, pushing it to develop a more granular understanding of nuanced trade-offs. This not only doubles the utility of our annotated tie data but also leads to a more stable training dynamic. As illustrated in Appendix A.7, our tie-disentanglement method results in a smoother training loss curve and more consistent performance gains on the validation set. Figure 3 shows some examples of our reward model giving rewards that are aligned with humans.
@@ 4.1 Implementation Details
We train our reward model, EditReward, using 200K high-quality pairwise preference samples from our dataset. For our main experimental results, we report performance using two powerful vision-language models as backbones: Qwen2.5-VL-7B and MiMo-VL-7B. To ensure a controlled comparison, all ablation studies are conducted consistently using the Qwen2.5-VL-7B backbone. During training, all parameters of the backbone are unfrozen and set as trainable. The training is performed for 2 epochs on a cluster of 8 NVIDIA A100 GPUs. For preprocessing, all training images are resized to $448\times 448$ pixels while preserving their original aspect ratios. Additional training details are provided in Appendix A.4.
@@ 4.2 Benchmarks and Baselines
Existing Benchmarks. We utilize ImagenHub (Ku et al., 2024), GenAI-Bench (Jiang et al., 2024), and AURORA-Bench (Krojer et al., 2024). For benchmarks with point-wise annotations like ImagenHub, we measure the Spearman rank correlation to assess alignment with human scores. For ImagenHub, which includes three ratings per sample, we also compute the Human-to-Human correlation as a practical upper bound (Ku et al., 2023). For benchmarks with paired comparisons like GenAI-Bench and the pair-wise split of AURORA-Bench, we report the prediction accuracy. Additional details of the evaluation across different methods are provided in Appendix A.5.
@@ 4.4 Application: EditReward as a Reward
To demonstrate EditReward’s practical utility as a data supervisor, we conducted a data curation experiment designed to improve a state-of-the-art editing model. We employed our reward model to score the  46,000 examples in the ShareGPT-4o-Image dataset (Chen et al., 2025a), from which we selected a high-quality subset of the top 20,000 samples. This curated dataset was then used to fine-tune the powerful Step1X-Edit model (Liu et al., 2025b).
Table 3: Comprehensive comparison of state-of-the-art models on both the English and Chinese versions of the GEdit-Bench benchmark, across intersection and full test sets. Our model, significantly improve the base model Step1X-Edit. $\uparrow$‘ indicates higher the better. *-I means intersection set.
@@
Ours (EditReward as reward) 7.895 6.946 7.131 7.854 6.931 7.086 7.757 7.024 7.074 7.658 6.995 7.001
Evaluation Protocol. To measure the impact of this curation, we evaluate the resulting model on the comprehensive GEdit-Bench. This benchmark features both English (EN) and Chinese (CN) instructions, as well as a challenging ”Intersection” subset containing prompts that all models could process. We compare our fine-tuned model against a diverse range of baselines, including the original Step1X-Edit, the same model fine-tuned on the full unfiltered dataset, and other leading open-source and proprietary models like Doubao and GPT-Image-1. Following established practices Ku et al. (2023), performance is judged by GPT-4o on three metrics (0-10 scale): Semantic Consistency (G_SQ) for instruction fidelity, Perceptual Quality (G_PQ) for visual realism, and an Overall Score (G_O) for overall quality.
Results and Analysis. As detailed in Table 3, this reward-driven filtering yields significant performance gains. On the GEdit-Bench Overall score, our fine-tuned model achieves 7.086, substantially outperforming both the original Step1X-Edit baseline (6.444) and the model trained on the full, unfiltered dataset (6.780). This finding is crucial, as it confirms that data quality, as judged by our reward model, is more impactful than sheer data quantity. EditReward successfully prunes noisy examples that would otherwise degrade performance during fine-tuning. This uplift elevates the open-source Step1X-Edit to be competitive with top-tier editors like Doubao, validating our model’s potential as an essential tool for training next-generation generative models.
@@ 5 Related Works
Evolution of Instruction-Guided Image Editing. Instruction-guided image editing has rapidly evolved from early trajectory-based methods. Diffusion models (Song et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2022; Podell et al., 2023) first enabled editing via dual-prompt formulations that relied on cross-attention manipulation or inversion (Hertz et al., 2022; Mokady et al., 2023; Wallace et al., 2023). The paradigm then shifted to more user-friendly single-instruction editing, pioneered by InstructPix2Pix (Brooks et al., 2023) and refined by works like MagicBrush and Emu-Edit (Zhang et al., 2024a; b; Sheynin et al., 2024) that focused on curating high-quality datasets. This trajectory-based family has been further advanced by flow-matching models (BlackForestLabs et al., 2025), which improve training and sampling efficiency. In parallel, sequential generative models, including autoregressive approaches (Yu et al., 2022; Tian et al., 2024), enhance compositional reasoning. The most recent advances feature hybrid multimodal architectures like OmniGen2 (Wu et al., 2025b) and BAGEL (Deng et al., 2025), which integrate large vision–language backbones with generative decoders to enable more context-aware, conversational editing.
Evaluating Instruction-Guided Image Editing. Early evaluation of image editing relied on perceptual metrics like LPIPS (Zhang et al., 2018), but these require reference images and fail to assess semantic alignment. CLIP-based metrics (Hessel et al., 2021) were introduced for text–image consistency but also show limited correlation with human judgment (Ku et al., 2024). The advent of large vision–language models (VLMs) enabled zero-shot evaluation, with proprietary models (Ku et al., 2023; Wang et al., 2025b) demonstrating promising human correlation while open-source counterparts (Liu et al., 2023; Laurençon et al., 2024) have lagged (Jiang et al., 2024). Consequently, recent work has focused on improving open-source evaluators via fine-tuning. One strategy distills supervision from proprietary models (Wei et al., 2024; Gu et al., 2024; Wu et al., 2024), which risks inheriting model biases. The other collects direct human annotations (Xu et al., 2024; Liang et al., 2024; Wu et al., 2023), offering higher-quality signals but typically at a smaller scale. Our work contributes a large-scale, expert-annotated dataset, enabling more reliable and robust reward modeling for image editing.
@@ 6 Conclusion
In this paper, we addressed the critical bottleneck hindering the advancement of open-source instruction-guided image editing: the lack of a reliable, human-aligned reward model for scaling up high-quality training data. To this end, we introduced a three-part solution: (1) EditReward-Data, a new large-scale (200K) preference dataset curated with rigorous expert annotation to minimize the noise and bias prevalent in existing resources; (2) EditReward, a dedicated reward model trained on this high-fidelity data to specialize in the image editing domain; and (3) EditReward-Bench, a challenging new benchmark featuring multi-way preference tasks to enable more robust evaluation. Our experimental results validate the effectiveness of our approach. EditReward establishes a new state of the art, demonstrating superior correlation with human judgment by outperforming strong VLM judges like GPT-5 and GPT-4o on public benchmarks. More importantly, we demonstrated its practical utility in a downstream data curation task: fine-tuning Step1X-Edit on a 20K subset of data filtered by EditReward yielded significantly better performance than training on the full 46K noisy dataset (7.1 vs. 6.7 overall score on GEdit-Bench). This confirms that a high-quality reward signal is a key ingredient for training powerful, next-generation editing models. Ultimately, this work provides both a methodology and a set of open resources to help bridge the gap between open-source and proprietary image editing models. This strong alignment with human judgment confirms EditReward’s capability as a data supervisor and highlights its potential for more advanced applications, including reinforcement learning-based post-training and test-time scaling of image editing models. To empower the community and facilitate future research, we will publicly release our EditReward-Data dataset, the trained EditReward model, and the EditReward-Bench benchmark.
@@
The development of advanced instruction-guided image editing models, which our work aims to evaluate and improve, carries significant ethical implications. While these technologies enable powerful creative expression, they can also be misused to generate deceptive or harmful content, such as deepfakes, misinformation, or fraudulent documents, lowering the barrier for malicious actors. Our work, by creating a more effective reward model, could inadvertently contribute to accelerating these capabilities. We acknowledge this dual-use potential and have taken steps to mitigate risks. The EditReward-Data dataset was constructed from publicly available, non-sensitive benchmarks, and we have made efforts to filter out any potentially harmful or personally identifiable content. Our reward model, EditReward, is trained to align with constructive and high-quality edits, as defined by our multi-dimensional rubric, not harmful instructions. By publicly releasing our dataset, model, and code, we aim to promote transparency and enable the research community to further study the safety, biases, and alignment of such models.

### v2

# arXiv:2509.26346v2[cs.CV] 28 Feb 2026

@@ Abstract
Recently, we have witnessed great progress in image editing with natural language instructions. Several closed-source models like GPT-Image-1, Seedream, and Google-Nano-Banana have shown highly promising progress. However, the open-source models are still lagging. The main bottleneck is the lack of a reliable reward model to scale up high-quality synthetic training data. To address this critical bottleneck, we built EditReward, trained with our new large-scale human preference dataset, meticulously annotated by trained experts following a rigorous protocol containing over 200K preference pairs. EditReward demonstrates superior alignment with human preferences in instruction-guided image editing tasks. Experiments show that EditReward achieves state-of-the-art human correlation on established benchmarks such as GenAI-Bench, AURORA-Bench, ImagenHub, and our new EditReward-Bench, outperforming a wide range of VLM-as-judge models. Furthermore, we use EditReward to select a high-quality subset from the existing noisy ShareGPT-4o-Image dataset. We train Step1X-Edit on the selected subset, which shows significant improvement over training on the full set. This demonstrates EditReward’s ability to serve as a reward model to scale up high-quality training data for image editing. EditReward with its training dataset will be released to help the community build more high-quality image editing training datasets to catch up with the frontier ones.
@@ 2.1 The EditReward-Data Construction
EditReward-Data contains 9557 instruction–image pairs collected from six established editing benchmarks: GEdit-Bench (606) (Liu et al., 2025b), ImgEdit-Bench (737) (Ye et al., 2025), MagicBrush (1,053) (Zhang et al., 2024a), AnyEdit (1,250) (Yu et al., 2025), EmuEdit (5,611) (Sheynin et al., 2024), and an internal set (300). This aggregation ensures broad coverage of semantically grounded and executable editing instructions. For each instruction, we generated 12 candidate images using six state-of-the-art models: Step1X-Edit (Liu et al., 2025b), Flux-Kontext (BlackForestLabs et al., 2025), Qwen-Image-Edit (Wu et al., 2025a), BAGEL (Deng et al., 2025), Ovis-U1 (Wang et al., 2025a), and OmniGen2 (Wu et al., 2025b), with multiple random seeds to avoid model bias. Seven candidates were randomly sampled for human evaluation. Annotators scored each image on a 4-point Likert scale (1 = Poor and 4 = Excellent) along two dimensions: Instruction Following (semantic accuracy, completeness, and no unprompted changes) and Visual Quality (plausibility, artifact-free rendering, and aesthetics). This rubric yields more informative labels than single-score schemes. Details of the annotation protocol and quality-control process are provided in the Appendix  A.2. Comprehensive statistics of the dataset are provided in Table 1 and Figure 2. EditReward-Data is unique in combining large-scale, expert human annotation and a multi-dimensional scoring rubric, making it a strong foundation for training editing reward models. More representative examples from EditReward-Data are shown in the Appendix  A.10. The IAA results in Table 9 provide a critical quantitative assessment of our expert annotation quality (Fleiss, 1971). We highlight the values derived from Krippendorff’s Alpha ($\alpha$(Krippendorff, 2011), which is the most appropriate metric as it correctly models the ordinal nature of our 4-point Likert scale. The $\alpha$ scores of $\mathbf{0.668}$ for Instruction Following (IF) and $\mathbf{0.597}$ for Visual Quality (VQ) establish a strong, quantified baseline for human consistency. Crucially, the observed difference ($\text{IAA}_{\text{IF}}>\text{IAA}_{\text{VQ}}$) provides empirical validation for our core contribution: it confirms that the VQ dimension is inherently more subjective than IF. This validates our design choice to use a multi-dimensional rubric and a multi-head reward model, as a single holistic score would obscure this critical difference in human variance.
@@
Table 2: Inter-Annotator Agreement (IAA) Metrics for Expert Annotations. The robust Krippendorff Alpha ($\alpha$) values confirm the high reliability of our multi-dimensional expert scoring.
Dataset Fleiss’ Kappa (IF) Fleiss’ Kappa (VQ) Krippendorff Alpha (IF) Krippendorff Alpha (VQ)
EditReward-Data 0.4157 0.3203 0.6762 0.5720
EditReward-Bench 0.3962 0.3157 0.6623 0.6114
All data 0.3994 0.3111 0.6685 0.5972
Figure 2: Statistics of our EditReward-Data and EditReward-Bench.
@@ 3.2 Multi-Dimensional Uncertainty-Aware Ranking
This strategy forces the model to reconcile seemingly contradictory signals for the same input pair, pushing it to develop a more granular understanding of nuanced trade-offs. This not only doubles the utility of our annotated tie data but also leads to a more stable training dynamic. As illustrated in Appendix A.7, our tie-disentanglement method results in a smoother training loss curve and more consistent performance gains on the validation set. Figure 3 shows some examples of our reward model giving rewards that are aligned with humans. More failure mode anaylysis of EditReward is shown in th Appendix  A.13.
@@ 4.1 Implementation Details
We train our reward model, EditReward, using 200K high-quality pairwise preference samples from our dataset. For our main experimental results, we report performance using two powerful vision-language models as backbones: Qwen2.5-VL-7B and MiMo-VL-7B. To ensure a controlled comparison, all ablation studies are conducted consistently using the Qwen2.5-VL-7B backbone. During training, all parameters of the backbone are unfrozen and set as trainable. The training is performed for 2 epochs on a cluster of 8 NVIDIA A800 GPUs. We follow the hyperparameter configuration to HPSv3, using a learning rate of $2\times 10^{-6}$ with a cosine learning rate schedule and a warm-up ratio of 0.05. With a per-GPU batch size of 2, the total effective batch size is 16. For preprocessing, all training images are resized to $448\times 448$ pixels while preserving their original aspect ratios. Additional training details are provided in Appendix A.4.
@@ 4.2 Benchmarks and Baselines
Existing Benchmarks. We utilize ImagenHub (Ku et al., 2024), GenAI-Bench (Jiang et al., 2024), and AURORA-Bench (Krojer et al., 2024). We explicitly confirm that we verified no overlap exists between our EditReward-Data training set and these evaluation benchmarks. They serve as fully independent, held-out testbeds, ensuring a fair and unbiased evaluation of EditReward’s generalization. For benchmarks with point-wise annotations like ImagenHub, we measure the Spearman rank correlation to assess alignment with human scores. For ImagenHub, which includes three ratings per sample, we also compute the Human-to-Human correlation as a practical upper bound (Ku et al., 2023). For benchmarks with paired comparisons like GenAI-Bench and the pair-wise split of AURORA-Bench, we report the prediction accuracy. Additional details of the evaluation across different methods are provided in Appendix A.5.
@@ 4.4 Application: EditReward as a Reward
To demonstrate EditReward’s practical utility as a data supervisor, we conducted a data curation experiment designed to improve a state-of-the-art editing model. We employed our reward model to score the  46,000 examples in the ShareGPT-4o-Image dataset (Chen et al., 2025a), from which we selected high-quality subsets (Top 10K, 20K, and 30K) for comparative analysis. This curated dataset was then used to fine-tune the powerful Step1X-Edit model (Liu et al., 2025b). The computational cost for scoring the samples in this pool was minimal, requiring only average 2.61 GPU hours, demonstrating high efficiency (0.25 seconds/sample).
Table 4: Comprehensive comparison of state-of-the-art models on both the English and Chinese versions of the GEdit-Bench benchmark, across intersection and full test sets. Our model, significantly improve the base model Step1X-Edit, including sensitivity analysis for the EditReward-curated subsets (Top 10K, 20K, and 30K). $\uparrow$ indicates higher the better. *-I means intersection set.
@@
Ours (EditReward as reward) (Top-K Sensitivity)
Step1X-Edit + Ours (Top 10K) 7.762 6.811 6.957 7.690 6.866 6.938 7.591 7.064 7.000 7.591 7.047 6.987
Step1X-Edit + Ours (Top 30K) 7.641 6.957 7.007 7.632 6.890 6.962 7.524 7.068 6.938 7.456 7.098 6.888
Step1X-Edit + Ours (Top 20K) 7.895 6.946 7.131 7.854 6.931 7.086 7.757 7.024 7.074 7.658 6.995 7.001
Evaluation Protocol. To measure the impact of this curation, we evaluate the resulting model on the comprehensive GEdit-Bench. This benchmark features both English (EN) and Chinese (CN) instructions, as well as a challenging ”Intersection” subset containing prompts that all models could process. We compare our fine-tuned model against a diverse range of baselines, including the original Step1X-Edit, the same model fine-tuned on the full unfiltered dataset, and other leading open-source and proprietary models like Doubao and GPT-Image-1. Following established practices Ku et al. (2023), performance is judged by GPT-4o on three metrics (0-10 scale): Semantic Consistency (G_SC) for instruction fidelity, Perceptual Quality (G_PQ) for visual realism, and an Overall Score (G_O) for overall quality.
Results and Analysis. As detailed in Table 4, this reward-driven filtering yields significant performance gains. The results show a clear trade-off between data quality and volume, confirming that our EditReward-filtered 20K subset represents the optimal balance for fine-tuning. Our best-performing model, trained on the Top 20K subset, achieves an English G_O score of 7.086. This substantially outperforms the original Step1X-Edit baseline (6.444) and the model trained on the full, noisy 46K dataset (6.780). Furthermore, our sensitivity analysis confirms that while the Top 10K subset (representing the highest signal-to-noise ratio) also outperforms the full set (G_O: 6.938), it is marginally inferior to the Top 20K subset, indicating the 20K size is necessary for robust generalization and avoiding underfitting. Crucially, the Top 30K subset (G_O: 6.962) yields diminishing returns compared to the Top 20K, confirming that including lower-quality data dilutes the training signal. This finding is crucial, as it confirms that data quality, as judged by our reward model, is more impactful than sheer data quantity. EditReward successfully prunes noisy examples that would otherwise degrade performance during fine-tuning. This uplift elevates the open-source Step1X-Edit to be competitive with top-tier editors like Doubao, validating our model’s potential as an essential tool for training next-generation generative models.
Table 5: Accuracy comparison on OOD tasks (Text & Style) sourced from Open Images. EditReward achieves performance comparable to GPT-4o while being open-source and cost-effective.
Model Text Category Style Category Overall
GPT-4o 45.50 35.79 41.69
EditReward (on MiMo-VL-7B-SFT) 47.83 45.41 46.80
4.5 Out-of-Distribution Generalization Analysis
To evaluate robustness outside the training pool, we conducted a targeted experiment on two challenging Out-of-Distribution (OOD) categories: Text-in-Image (OCR) and Style Transfer.
Experimental Setup. We constructed a specialized OOD set sourced from Open Images (distinct from training sources), comprising 253 Text pairs and 185 Style pairs with expert annotations. We compare EditReward (on MiMo-VL-7B-SFT) against the commercial SOTA, GPT-4o.
Results. Table 5 shows EditReward achieves performance comparable to GPT-4o on these tasks, maintaining competitive alignment despite inherent VLM difficulties with OCR. Crucially, EditReward offers significant advantages as a cost-effective, open-source alternative with faster inference speeds.
@@ 5 Related Works
Evolution of Instruction-Guided Image Editing. Instruction-guided image editing has rapidly evolved from early trajectory-based methods. Diffusion models (Song et al., 2020; Dhariwal and Nichol, 2021; Rombach et al., 2022; Podell et al., 2023) first enabled editing via dual-prompt formulations that relied on cross-attention manipulation or inversion (Hertz et al., 2022; Mokady et al., 2023; Wallace et al., 2023). The paradigm then shifted to more user-friendly single-instruction editing, pioneered by InstructPix2Pix (Brooks et al., 2023) and refined by works like MagicBrush and Emu-Edit (Zhang et al., 2024a; b; Sheynin et al., 2024) that focused on curating high-quality datasets. This trajectory-based family has been further advanced by flow-matching models (BlackForestLabs et al., 2025), which improve training and sampling efficiency. In parallel, sequential generative models, including autoregressive approaches (Yu et al., 2022; Tian et al., 2024), enhance compositional reasoning. The most recent advances feature hybrid multimodal architectures like OmniGen2 (Wu et al., 2025b) and BAGEL (Deng et al., 2025), which integrate large vision–language backbones with generative decoders to enable more context-aware, conversational editing.
Evaluating Instruction-Guided Image Editing. Early evaluation of image editing relied on perceptual metrics like LPIPS (Zhang et al., 2018), but these require reference images and fail to assess semantic alignment. CLIP-based metrics (Hessel et al., 2021) were introduced for text–image consistency but also show limited correlation with human judgment (Ku et al., 2024). The advent of large vision–language models (VLMs) enabled zero-shot evaluation, with proprietary models (Ku et al., 2023; Wang et al., 2025b) demonstrating promising human correlation while open-source counterparts (Liu et al., 2023; Laurençon et al., 2024) have lagged (Jiang et al., 2024). Consequently, recent work has focused on improving open-source evaluators via fine-tuning. One strategy distills supervision from proprietary models (Wei et al., 2024; Gu et al., 2024; Wu et al., 2024), which risks inheriting model biases. The other collects direct human annotations (Xu et al., 2024; Liang et al., 2024; Wu et al., 2023; Sani et al., 2026), offering higher-quality signals but typically at a smaller scale. Our work contributes a large-scale, expert-annotated dataset, enabling more reliable and robust reward modeling for image editing.
@@ 6 Conclusion
In this paper, we addressed the critical bottleneck hindering the advancement of open-source instruction-guided image editing: the lack of a reliable, human-aligned reward model for scaling up high-quality training data. To this end, we introduced a three-part solution: (1) EditReward-Data, a new large-scale (200K) preference dataset curated with rigorous expert annotation to minimize the noise and bias prevalent in existing resources; (2) EditReward, a dedicated reward model trained on this high-fidelity data to specialize in the image editing domain; and (3) EditReward-Bench, a challenging new benchmark featuring multi-way preference tasks to enable more robust evaluation. Our experimental results validate the effectiveness of our approach. EditReward establishes a new state of the art, demonstrating superior correlation with human judgment by outperforming strong VLM judges like GPT-5 and GPT-4o on public benchmarks. More importantly, we demonstrated its practical utility in a downstream data curation task: fine-tuning Step1X-Edit on a 20K subset of data filtered by EditReward yielded significantly better performance than training on the full 46K noisy dataset (7.1 vs. 6.7 overall score on GEdit-Bench). This confirms that a high-quality reward signal is a key ingredient for training powerful, next-generation editing models. Ultimately, this work provides both a methodology and a set of open resources to help bridge the gap between open-source and proprietary image editing models. To empower the community and facilitate future research, we will publicly release our EditReward-Data dataset, the trained EditReward model, and the EditReward-Bench benchmark.
@@
The development of advanced instruction-guided image editing models, which our work aims to evaluate and improve, carries significant ethical implications. While these technologies enable powerful creative expression, they can also be misused to generate deceptive or harmful content, such as deepfakes, misinformation, or fraudulent documents, lowering the barrier for malicious actors. Our work, by creating a more effective reward model, could inadvertently contribute to accelerating these capabilities. We acknowledge this dual-use potential and have taken steps to mitigate risks. Specifically, the EditReward-Data dataset was constructed from publicly available, non-sensitive benchmarks, and automated and manual filtering was applied to remove any personally identifiable information (PII) or sensitive content. Our reward model, EditReward, is trained to align with constructive and high-quality edits, as defined by our multi-dimensional rubric, and does not follow harmful or malicious instructions. Additionally, all generated data and model outputs will be released under a CC-BY-NC-SA 4.0 license, explicitly prohibiting commercial use, which mitigates potential misuse such as the creation of deepfakes or other harmful applications. By publicly releasing our dataset, model, and code, we aim to promote transparency and enable the research community to further study the safety, biases, and alignment of such models. Finally, we encourage the community to adopt similar safeguards, including watermarking, provenance tracking, and careful curation of training data, when deploying or extending instruction-guided image editing models.